# Non-IID data in Federated Learning: A Systematic Review with Taxonomy, Metrics, Methods, Frameworks and Future Directions

### v1

@@ X Conclusion
In summary, this work offers a comprehensive systematic review of the state-of-the-art non-IID data in FL. Unlike existing surveys and reviews, we include a detailed taxonomy for non-IID data, partition protocols and metrics to quantify non-IIDness, popular solutions to address data heterogeneity, and standardized frameworks employed in FL with heterogeneous data. We include modality skew as a novel category in the taxonomy data heterogeneity, which appears as a new trend among the researchers. We indicate a lack of consensus on non-IID classification and quantification, highlighting the importance of using metrics to measure non-IIDness. We claim that quantifying only one skew type at a time is insufficient for fully understanding and addressing non-IIDness. Finally, we suggest some research directions that can be adopted in future investigations in the field.

### v2

@@ X Conclusion
In summary, this work offers a comprehensive technical survey of the state-of-the-art non-IID data in FL. Unlike existing surveys, we include a detailed taxonomy for non-IID data, partition protocols and metrics to quantify non-IIDness, popular solutions to address data heterogeneity, and standardized frameworks employed in FL with heterogeneous data. We include modality skew as a novel category in the taxonomy data heterogeneity, which appears as a new trend among the researchers. We indicate a lack of consensus on non-IID classification and quantification, highlighting the importance of using metrics to measure non-IIDness. We claim that quantifying only one skew type at a time is insufficient for fully understanding and addressing non-IIDness. Finally, we suggest some research directions that can be adopted in future investigations in the field.